OpenStudy (anonymous):

2x+5y=70 into y=mx+b form

OpenStudy (anonymous):

Do you have an answer choice?

OpenStudy (ranga):

subtract 2x from both sides. Then divide EVERY term by 5. That includes dividing the 5y, the -2x and the 70.

OpenStudy (anonymous):

2x+5y=70 -2x -2x Subtract 2x from each side ------------------ 5y=-2x+70 Divide each side by 5 ----- --------- 5 5 y= -2/5+14